Three consecutive integers are such that the square of the third is 76
more thant the square of the second. Find the integers.

Answer by checkley71(8403)

THE THREE COSECUTIVE INTEGERS ARE X, X+1 & X+2.
(X+2)^2=(X+1)^2+76
X^2+4X+4=X^2+2X+1+76
4X-2X=-4+1+76
2X=73
X=73/2
X=36.5
X+1=37.5 ANSWER FOR THE SECOND NUMBER
X=2=38.5 ANSWER FOR THE THIRD NUMBER.
PROOF
(38.5^2=37.5^2+76
1482.25=1406.25+76
1482.25=1482.25